2013-APR-26: MONET/North derotator (3rd telescope axis) is acting up - the system remains unavailable until we can figure out what's wrong.
2013-MAR-19: MONET/North returns to non-normal operation (use only with an admin in the background).
2013-MAR-18: MONET/North roof having synchronisation problems - the system remains unavailable until McD Physical Plant checks things and gives their OK
2013-JAN-22: MONET/North returns to non-normal operation (use only with an admin in the background).
2013-JAN-16: MONET/North remains closed due to power-line instabilities which are causing the UPS to cycle constantly.
2013-JAN-13: MONET/North closed due to extremely cold weather and snow.
2012-DEC-29: MONET/North can be operated using the present UPS batteries.
2012-DEC-26: MONET/North shutdown due to faulty UPS battery; unknown how long it will take to replace.
2012-NOV-12: MONET/North total shutdown due to accidental capping of the electricity to the mountain by a construction crew.
2012-JUL-24: The MONET/North telescope will be down for maintenance from 26 JUL to 4 AUG (removal and cleaning of the mirrors).
2012-FEB-01: MONET/North down for computer replacements for 1-2 days.
2011-DEC-23: McD PP says roofs are OK now.
2011-DEC-22: because of a roof problem, the observatory is closed until McD PP says things are fine again.
2011-NOV-28: the telescope computer died and has to be replaced - will take a few days....
2011-OCT: REMEMBER that operations are possible only if you have an admin available in the background!
2010-AUG-01: Due to a sudden storm MONET/North got rained on and cannot be operated without an accompanying admin until further notice.*
2011-JUL-28: after a couple of power cycles, the telescope electronics decided that things were fine after all. Back to work!
2011-JUL-27: a day after re-release of the telescope, there is an azimuth controller problem which refuses to clear: we're working on it.....
2011-MAY-23: the MONET/North telescope will be shut down until the end of June for a major roof overhaul.
2011-APR-21: PLEASE WATCH FOR DUST - operation is not allowed for values over 60,000 ppcf, so the portal may be blocked on particularly dusty days (red light in control room).
2011-APR-14: due to heavy residual smoke and dust, we'll have to wait another day or so for normal operations to resume.
2011-APR-11: Here a report from C. Wilkenson/Univ. Texas: There was a wildfire in Fort Davis Saturday which knocked out a bunch of utility poles, burned some homes, and forced evacuations throughout the area. It started in Marfa, but high winds drove it through Fort Davis and toward Balmorhea. That fire is still massive and out of control, but at least it is not threatening the Observatory at this point. Unfortunately, we are without power and it is uncertain when that will be resolved. We have heard dates ranging from Monday to Friday. The site is running on minimal staff, as water is running low (electric pumps + no power = minimal water). I suspect there will be several days of "restart" before the site is running at 100% again.
2011-APR-06: camera daemon causing problems.....
2011-APR-06: the West Texas network is back, the daemons restarted, all ready to go.
2011-APR-05: massive network failure at McDonald of unknown severity.
2011-MAR-29: TT tests will be performed every day after 12:00 UT.
2011-MAR-5: due to quick work by McDonald staff, the old motor controller is back and the telescope working (for now).
2011-MAR-4: the motor controller is out again; the telescope will be down until this problem is fixed.
2011-MAR-01: the network card on the telescope computer had to be replaced.
2011-FEB-28: the motor controller for the azimuth axis has been replaced but the system still has glitches - expect the telescope to be on and off for the next few days until we find out what's happening.
2010-DEC-02: the network connection to McDonald has failed completely, presumedly a local problem (e.g. switch).
2010-SEP-04: New pointing model installed but not tested. The roof did not close yesterday when the humidity rose, so observers MUST monitor the weather faithfully (if problem isn't fixed soon, we'll have to close for safety reasons). The "EXPOSE" button hangs up after turning into "STOP": a window refresh "fixes" this old problem returned to haunt us.
2010-AUG-23: Warning - the current pointing model is performing so badly, that the tracking has become totally unacceptable. We will be trying to update the pointing model within the next days. Until then, you'll want to use very short exposures.
2010-AUG-19: Normal use of MONET/North starting Monday, the 23rd of August!
2010-JUL-04: A few weeks turned into a few months. MONET/North has been re-instated while the final roof fix is performed at MONET/South (her northern sister-telescope follows a month or so later), but is being tested for a few more days (upgrades, new pointing model). Stay posted!
2010-FEB-24: MONET/North will be shut down for about two weeks in order to investigate the roof mechanism (in the hopes of finding the bug in the MONET/South roof).
2009-JAN-11: Note that the weather info shown in the telescope operation GUI has occasionally not been correct - we are looking for the bug, so please keep a good eye on the other weather stations.
2010-JAN-11: Due to a faulty firmware update by the manufacturers, the roof on MONET/North cannot be operated until further notice.
2009-DEC-11: The roof has been left in a "technician mode" by the local engineer due to mechanical tests and so cannot be opened until the tests are complete.
2009-DEC-07: The upgrade of the MONET/North roof control system is complete, but the final testing couldn't be completed due to heavy winds and dust, so expect interruptions.
2009-OCT-14: the filter wheel cables have been fixed - thanks to the efforts of the McDonald? staff - so normal usage has been restored.* (Hessman)
Rules of Conduct:
Please remember to enter your observations in the observing calendar, even if they are planned at the last moment.
If something goes wrong, make sure you also notify the observer after you (you should have received a list of current observers)!
Enter a log report for/after ever use of the telescopes, especially if there are problems; it doesn't have to be fancy - just some simple sentences saying what you did and how it went.